Find all the zeroes for f(x)=x^3+4x^2-5x and write the function f(x) as a product of linear fsctors. I know how to find the real zeroes using the Rational Zeroes Test but I think I need to find the complex zeroes as well. I don't know how to find the complex zeroes without being given one complex zero because of the complex conjugate pair.
Answer by ewatrrr(24785)

 
Hi,
f(x)=x^3+4x^2-5x = x(x^2 + 4x - 5) = x(x+5)(x-1) = 0 ||| zeroes are: -5,0,1